# Resonant dynamics of spin cluster in a periodically driven one-dimensional Rydberg lattice

### v1

@@ Abstract
Rydberg lattice under facilitation conditions can feature kinetic constraints, leading to ballistic and non-ergodic behavior at different detuning intensities. Here, we demonstrate that a resonant driving field can achieve effects similar to those under facilitation conditions. We focus on the relaxation dynamics of spin clusters in a periodically driven Rydberg spin lattice. Through an effective Hamiltonian for the domain walls of the spin cluster, it is shown that when the driving frequency is resonant with the Rydberg interaction, the spin cluster exhibits ballistic expansion with half the spreading rate compared to the case of facilitation conditions. However, near the resonant point, the spin cluster displays confinement behavior of the Bloch-like oscillations. These results demonstrate the rich dynamic behaviors in the driven Rydberg spin lattices and may find applications in quantum state manipulation.
@@ I Introduction
Quantum systems subjected to external fields have garnered significant interest in recent years due to their potential to exhibit novel nonequilibrium phenomena and promising applications[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], such as Floquet topological phases [1], discrete time crystals [2, 3], dynamical localization [6, 7, 8, 9, 10] and the control of quantum states through coherent driving [11, 12, 13, 14]. Particularly, under the resonant condition where the frequency or strength of an applied field matches the energy spacing of the system, the interplay between the field and the system are enhanced, resulting in substantial changes in terms of long-range quantum correlations, quantum thermalization, and dynamic properties of the systems [20, 21, 22, 23, 24, 25, 13, 26, 27, 28]. This realm bridges fundamental questions about non-equilibrium quantum matters with applications in quantum simulation and quantum information.
@@
The Rydberg lattice systems [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 28, 45, 46] are the ideal platforms for studying the resonant dynamics of many-body quantum states, not only for the abilities to the fine-tuning of particle interactions and external fields, but also for the state of the art measurement techniques for the quantum states [34]. Experimentally, the time-dependent driving fields can be readily implemented through the Rydberg lasers with time-varying Rabi frequency [38]. Recently, it is shown that the Rydberg lattices under facilitation conditions, where the laser detuning cancels out the nearest-neighbor interaction of atoms, feature rich dynamic behaviors [31, 47, 37, 48, 49, 44] including Bloch oscillations of spin cluster [37], coherent spin-phonon scattering [44] and non-classical spin-phonon correlations [49]. It turns out that the facilitation condition can be considered as a resonant condition for the spin clusters dynamics. Therefore, a resonant driving field may play a similar role to the facilitation condition in the aspect of spin clusters dynamics in the Rydberg lattice. The interplay between the driving field and Rydberg interaction may give rise to dynamics of spin cluster different from that under facilitation condition.
@@
In this work, we investigate the dynamics of spin cluster in a periodically driven one-dimensional (1D) Rydberg lattice. Specifically, we focus on the relaxation dynamics of spin cluster in this system under high-frequency driving field with the form $\mathit{\Omega}\left(t\right)=\mathit{\Omega}_{0}\cos\left(\omega t\right)$. We derive an effective Hamiltonian for the domain walls of the spin cluster, and show that there is a resonant frequency. When the driving frequency $\omega$ is resonant with the Rydberg interaction $V_{i,j}$, the spin clusters exhibit ballistic expansion with half the spreading rate compared to the case of facilitation conditions before the domain walls of spin cluster collide. After the first collision, the spin clusters exhibit subdiffusive dynamics. While for an initially Gaussian distribution with momentum $\pi$, the spin cluster exhibits ballistic expansion, and the total Rydberg density is conserved in the entire time domain. These results are demonstrated by the time evolution of Rydberg density and its variance related to the mean square displacement. Moreover, we show that near the resonant driving frequency, the spin clusters manifest the coexistence of oscillation and expansion dynamics, and the Bloch-like oscillation of the spin cluster, for different initial excitations.
@@
The rest of this paper is organized as follows: In Sec. II, we introduce the Hamiltonian of the 1D Rydberg lattice with a periodic driving field, and discuss the dynamic properties in the spin cluster subspace under the condition of high frequency drive. In Sec. III, we derive the effective Hamiltonian for the domain walls of the spin cluster and obtain the resonant frequency. In Sec. IV, we investigate the dynamics in the resonant frequency and near the resonant frequency, and present numerical results of the evolutions of Rydberg density and its variance for different initial excitations. Finally, we conclude and discuss our findings in Sec. V.
@@ II Model and spin cluster subspace
where the operator $\hat{\sigma}_{i}^{x}=\left|\uparrow_{i}\right>\left<\downarrow_{i}\right|+\left|\downarrow_{i}\right>\left<\uparrow_{i}\right|$ couples the ground state $\left|\downarrow\right$> and the Rydberg state $\left|\uparrow\right$> of the $i$th atom through the periodic driving field with the form $\mathit{\Omega}\left(t\right)=\mathit{\Omega}_{0}\cos\left(\omega t\right)$, the operator $\hat{n}_{i}=\left|\uparrow_{i}\right>\left<\uparrow_{i}\right|$ projects onto the Rydberg state for the $i$th atom, and $V_{i,j}=\mathcal{V}_{0}/\left|i-j\right|^{6}$ is the long-range Van der Waals interaction between the Rydberg excited atoms at sites $i$ and $j$. Throughout this work, we assume $\mathit{\Omega}_{0}>0$ and $\mathcal{V}_{0}>0$, and periodic boundary conditions are applied. The dynamical controls for nonequilibrium quantum many-body states of this model can be implemented experimentally in neutral $^{87}\textrm{Rb}$ atoms trapped in optical tweezers [38]. In this work, we are interested in the dynamics of the spin cluster under a periodic driving field. To this end, we take the spin cluster excitation as an initial state, and analyze this problem under the Hamiltonian of driven Rydberg lattice in Eq. (1).
@@
with $0<j_{1}\leqslant j_{2}<N$, where the raising operator of the $j$th atom is defined as $\hat{\sigma}_{j}^{+}=\left|\uparrow_{j}\right>\left<\downarrow_{j}\right|$. Let us start with some Heuristic derivations. Firstly, noting that in this state, the ground-state atoms that are far away from the atoms inside the spin cluster, i.e., $l\ll j_{1}$ or $l\gg j_{2}$, are not subject to interactions $V_{ij}$. Thus before they are excited, they can be regarded as a series of isolated two-level atoms in a global driving field $\mathit{\Omega}\left(t\right)/2$, and each of them can be described by the Hamiltonian $\hat{H}_{l}(t)=\mathit{\Omega}\left(t\right)\hat{\sigma}_{l}^{x}/2$. Their transition to the Rydberg excited states will be suppressed by the driving field, provided that the driving frequency $\omega$ is high enough, namely $\omega\gg\mathit{\Omega}_{0}$. In fact, for an atom at the $l$th site that is initially prepared on the ground state $\left|\downarrow_{l}\right$>, it can be checked that after time $t$, the probability of transition to Rydberg state $\left|\uparrow_{l}\right$> is
@@
Figure 1: (a) Schematic illustration of the 1D Rydberg lattice under periodic driving field $\mathit{\Omega}\left(t\right)$. There are long-range interactions $V_{ij}$ between atoms in the Rydberg excited states $\left|\uparrow\right$>, and $\left|\downarrow\right$> represents the ground state of an atom. (b) and (c) present the numerical results of the Rydberg density $\left<\hat{n}_{j}\right$> for different initial excitations, under and near the resonant frequencies of driving fields, $\omega=5.0867$ and $\omega=5.2867$, respectively. Other parameters are taken as $N=20$, $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$.
Secondly, the interaction between the atoms in Rydberg states can be replaced by the effective detuning $\mathit{\Delta}_{\mathrm{eff}}\sim\mathcal{V}_{0}$, under the mean-field approximation. Then, inside the spin cluster, each atom can be described by the Hamiltonian $\hat{H}_{j}(t)=\mathit{\Omega}\left(t\right)\hat{\sigma}_{j}^{x}/2+\mathit{\Delta}_{\mathrm{eff}}\hat{n}_{j}$. According to the time-dependent perturbation theory, for an atom at site $j$ that is initially prepared in the Rydberg state $\left|\uparrow_{j}\right$>, the probability of transition to the ground state $\left|\downarrow_{j}\right$> at time $t$ is approximately
@@
which is suppressed when $\mathcal{V}_{0}\approx\omega\gg\mathit{\Omega}_{0}$ (see APPENDIX for the derivation). This achieves a similar effect to that under the facilitation conditions [31, 47, 37, 44, 49], where the detuning is set to cancel out the interaction between two adjacent atoms. In this parameter regime, it is interesting to further reveal the dynamic behavior of the domain walls of the spin cluster.
In order to gain some intuitions and preliminary conclusions about the spin cluster dynamics, we perform numerical computations of time evolutions, observing the expansion dynamics of the spin cluster when the driving frequency is taken as $\omega=5.0867$, and a slight deviation of the driving frequency leads to oscillation dynamics of the spin cluster, in Figs. 1(b) and (c), respectively. In Fig. 1(b), we present numerical results of the Rydberg density $\left<\hat{n}_{j}\right>(t)=\left<\Psi\left(t\right)\right|\hat{n}_{j}\left|\Psi\left(t\right)\right$> as a function of time for the initial excitation in Eq. (2), that is $\left|\Psi\left(0\right)\right>=\left|\psi\left(j_{1,}j_{2}\right)\right$> with $j_{1}=9$ and $j_{2}=12$, for the first case. The evolved state
@@
is calculated under the time-dependent Hamiltonian in Eq. (1) using the fourth-order Runge-Kutta method with time steps of length $\Delta t=4\times 10^{-4}$ and the Hamiltonian is updated in each time step, where $\hat{\mathcal{T}}$ is the time-ordering operator and $\mathrm{i}=\sqrt{-1}$. In Fig. 1(c), we take the a Gaussian distribution with width $\sigma=2$ of initial state as the following form
$\left|\Psi_{\mathrm{Gau.}}\left(0\right)\right>=\Lambda^{-1}\sum_{j_{1}}\mathrm{e}^{-\frac{\left(c-c_{0}\right)^{2}}{4\sigma^{2}}}\mathrm{e}^{-\mathrm{i}k_{0}c}\left|\psi\left(j_{1},j_{1}+r_{0}\right)\right$>, (6)
where $c=(j_{1}+j_{2})/2=j_{1}+(r_{0}-1)/2$ is the center of mass position of spin cluster (here $r_{0}=4$ is taken), $c_{0}=10.5$ and $k_{0}=\pi$ are the average center of mass position and momentum, and $\Lambda$ is the normalization constant. We analyze the mechanism of these dynamics in the next section.
@@ III Effective Hamiltonian and resonant frequency
which is a constant for large $r$. This means that for the fixed $k$ and $t$, the Hamiltonian $\hat{H}_{k}(t)$ describes a hopping particle in a lattice with skew potential $\mathcal{F}r$, which is closely related to the Bloch oscillations [50, 51].
In the large $N$ limit and for a finite $\mathcal{F}$, the eigenstates of the Hamiltonian in Eq. (9) are all localized, forming the celebrated Wannier-Stark ladder states [51]. We consider the eigenstates localized at $r\gg 1$, which have the form
@@
Figure 2: Time evolutions of Rydberg density $\left<\hat{n}_{j}\right$> and density variance $\delta\sigma(t)=\sigma(t)-\sigma(0)$ for the system under the resonant frequency of the driving field with different initial states. The frequency of the driving field is set as $\omega=5.0867$. The top panels present time evolution of Rydberg density $\left<\hat{n}_{j}\right$> for the initial excitations of (a1) simple spin cluster $\left|\psi\left(j_{1},j_{2}\right)\right$> with $j_{1}=46$ and $j_{2}=55$; the Gaussian distribution of the spin cluster in Eq. (6) with momentum (b1) $k_{0}=0$, (c1) $k_{0}=\pi/2$ and (d1) $k_{0}=\pi$. Other parameters for the Gaussian distribution are taken as $\sigma=4$, $c_{0}=50.5$ and $r_{0}=10$. The bottom panels (a2)-(d2) show the density variance $\delta\sigma(t)$ on double-logarithmic scales (black dots) corresponding to the Rydberg density in (a1)-(d1), and the solid lines represent the linear fittings for the data. Other parameters are taken as $N=100$, $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$.
Figure 3: Total Rydberg density $\mathcal{N}(t)$ as a function of time for the cases in Fig. 2(a1)-(d1). The frequency of the driving field is set as $\omega=5.0867$. Other parameters are taken as $N=100$, $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$.
@@
$\left|\Psi\left(0\right)\right>=\sum_{k}c_{k}\left|\Psi_{k}\left(0\right)\right>\otimes\left|k\right$>, (12)
@@
and left multiplied by $\left<\phi_{m,k}\right|$, we obtain the equation of motion for amplitude $\mathcal{C}_{n,k}(t)$, that is
@@
Note that, however, $\left|\phi_{n,k}\right\rangle$ is time-dependent. Submitting the expression of $\left|\phi_{n,k}\right\rangle$ [see Eq. (11)] into the above equation, and using the recurrence relation of Bessel function $2\partial_{x}J_{n}(x)=J_{n-1}(x)-J_{n+1}(x)$, we have

### v2

@@ Abstract
Rydberg lattice under facilitation conditions can feature kinetic constraints, leading to ballistic and nonergodic behavior at different detuning intensities. Here, we demonstrate that a resonant driving field can achieve effects similar to those under facilitation conditions. We focus on the relaxation dynamics of spin clusters in a periodically driven Rydberg spin lattice. Through an effective Hamiltonian for the domain walls of the spin cluster, it is shown that when the driving frequency is resonant with the Rydberg interaction, the spin cluster exhibits ballistic expansion with half the spreading rate compared to the case of facilitation conditions. However, near the resonant point, the spin cluster displays confinement behavior of the Bloch-like oscillations. These results demonstrate the rich dynamic behaviors in the driven Rydberg spin lattices and may find applications in quantum state manipulation.
@@ I Introduction
Quantum systems subjected to external fields have garnered significant interest in recent years due to their potential to exhibit novel nonequilibrium phenomena and promising applications[Rudner2020, else2020discrete, Collura2022, Longhi2012, Longhi2013, nag2014dynamical, luitz2017absence, agarwala2017effects, cao2022interaction, tiwari2024dynamical, bason2012high, Lin2014, zhang2022steady, mccullian2024coherent, li2017long, goldman2014periodically, zhang2024dynamics, zhang2024, zhang2024two], such as Floquet topological phases [Rudner2020], discrete time crystals [else2020discrete, Collura2022], dynamical localization [nag2014dynamical, luitz2017absence, agarwala2017effects, cao2022interaction, tiwari2024dynamical] and the control of quantum states through coherent driving [bason2012high, Lin2014, zhang2022steady, mccullian2024coherent]. Particularly, under the resonant condition where the frequency or strength of an applied field matches the energy spacing of the system, the interplay between the field and the system is enhanced, resulting in substantial changes in terms of long-range quantum correlations, quantum thermalization, and dynamic properties of the systems [SanchisAlepuz2007, goldman2015periodically, bhakuni2020drive, zhang2020resonant, haldar2021dynamical, xu2022coherent, zhang2022steady, li2024collective, liu2025resonant, giudici2025fast]. This realm bridges fundamental questions about nonequilibrium quantum matters with applications in quantum simulation and quantum information.
@@
The Rydberg lattice systems [Zeiher2016, Labuhn2016, Marcuzzi2017, Bernien2017, Yang2019, Browaeys2020, Mukherjee2020, Li2020, Magoni2021, Bluvstein2021, Scholl2021, ebadi2021quantum, Mukherjee2022, Anand2024, Chen2024, Magoni2024, giudici2025fast, Pal2025a, Bombieri2025] are the ideal platforms for studying the resonant dynamics of many-body quantum states, not only for the abilities for the fine-tuning of particle interactions and external fields, but also for the state of the art measurement techniques for the quantum states [Browaeys2020]. The realm of Rydberg lattice systems with time-dependent driving is a rapidly developing frontier that combines the strong Rydberg interactions, programmable lattice geometries, and Floquet engineering to provide a powerful platform for exploring and manipulating quantum many-body dynamics [Basak2018, guo2020optimized, zhao2023floquet, wu2025quantum]. Experimentally, the time-dependent driving fields can be readily implemented through the Rydberg lasers with time-varying Rabi frequency [Bluvstein2021]. Recently, it was shown that the Rydberg lattices under facilitation conditions, where the laser detuning cancels out the nearest-neighbor interaction of atoms, feature rich dynamic behaviors [Marcuzzi2017, Ostmann2019, Magoni2021, Magoni2022, Brady2025, Magoni2024] including Bloch oscillations of spin cluster [Magoni2021], coherent spin-phonon scattering [Magoni2024] and nonclassical spin-phonon correlations [Brady2025]. It turns out that the facilitation condition can be considered as a resonant condition for the spin cluster dynamics. Therefore, a resonant driving field may play a similar role in the facilitation condition in the aspect of spin cluster dynamics in the Rydberg lattice. The interplay between the driving field and Rydberg interaction may give rise to dynamics of spin cluster different from that under facilitation condition, and the modulated drive may provide another option or degree of freedom in parameter for manipulating the dynamics of spin clusters in the 1D Rydberg lattice.
@@
In this work, we investigate the dynamics of spin cluster in a periodically driven one-dimensional (1D) Rydberg lattice. Specifically, we focus on the relaxation dynamics of spin cluster in this system under high-frequency driving field with the form $\mathit{\Omega}\left(t\right)=\mathit{\Omega}_{0}\cos\left(\omega t\right)$. We derive an effective Hamiltonian for the domain walls of the spin cluster, and show that there is a resonant frequency. When the driving frequency $\omega$ is resonant with the Rydberg interaction $V_{i,j}$, the spin clusters exhibit ballistic expansion with half the spreading rate compared to the case of facilitation conditions before the domain walls of spin cluster collide. After the first collision, the spin clusters exhibit subdiffusive dynamics. While for an initially Gaussian distribution with momentum $\pi$, the spin cluster exhibits ballistic expansion, and the total Rydberg density is conserved in the entire time domain. These results are demonstrated by the time evolution of Rydberg density and its variance related to the mean-square displacement. Moreover, we show that near the resonant driving frequency, the spin clusters manifest the coexistence of oscillation and expansion dynamics, and the Bloch-like oscillation of the spin cluster, for different initial excitations.
@@
The rest of this paper is organized as follows: In Sec. II, we introduce the Hamiltonian of the 1D Rydberg lattice with a periodic driving field, and discuss the dynamic properties in the spin cluster subspace under the condition of high frequency drive. In Sec. III, we derive the effective Hamiltonian for the domain walls of the spin cluster and obtain the resonant frequency. In Sec. IV, we investigate the dynamics in the resonant frequency and near the resonant frequency, and present numerical results of the evolutions of Rydberg density and its variance for different initial excitations. Finally, we summarize and discuss our findings in Sec. V, and conclude the results in Sec. VI.
@@ II Model and spin cluster subspace
where the operator $\hat{\sigma}_{i}^{x}=\left|\uparrow_{i}\right>\left<\downarrow_{i}\right|+\left|\downarrow_{i}\right>\left<\uparrow_{i}\right|$ couples the ground state $\left|\downarrow\right$> and the Rydberg state $\left|\uparrow\right$> of the $i$th atom through the periodic driving field with the form $\mathit{\Omega}\left(t\right)=\mathit{\Omega}_{0}\cos\left(\omega t\right)$, the operator $\hat{n}_{i}=\left|\uparrow_{i}\right>\left<\uparrow_{i}\right|$ projects onto the Rydberg state for the $i$th atom, and $V_{i,j}=\mathcal{V}_{0}/\left|i-j\right|^{6}$ is the long-range Van der Waals interaction between the Rydberg excited atoms at sites $i$ and $j$. Throughout this work, we assume $\mathit{\Omega}_{0}>0$ and $\mathcal{V}_{0}>0$, and periodic boundary conditions are applied. The dynamical controls for nonequilibrium quantum many-body states of this model can be implemented experimentally in neutral $^{87}\textrm{Rb}$ atoms trapped in optical tweezers [Bluvstein2021]. In this work, we are interested in the dynamics of the spin cluster under a periodic driving field. To this end, we take the spin cluster excitation as an initial state, and analyze this problem under the Hamiltonian of driven Rydberg lattice in Eq. (1).
@@
with $0<j_{1}\leqslant j_{2}<N$, where the raising operator of the $j$th atom is defined as $\hat{\sigma}_{j}^{+}=\left|\uparrow_{j}\right>\left<\downarrow_{j}\right|$. Let us start with some heuristic derivations. First, note that in this state, the ground-state atoms that are far away from the atoms inside the spin cluster, i.e., $l\ll j_{1}$ or $l\gg j_{2}$, are not subject to interactions $V_{ij}$. Thus before they are excited, they can be regarded as a series of isolated two-level atoms in a global driving field $\mathit{\Omega}\left(t\right)/2$, and each of them can be described by the Hamiltonian $\hat{H}_{l}(t)=\mathit{\Omega}\left(t\right)\hat{\sigma}_{l}^{x}/2$. Their transition to the Rydberg excited states will be suppressed by the driving field, provided that the driving frequency $\omega$ is high enough, namely $\omega\gg\mathit{\Omega}_{0}$. In fact, for an atom at the $l$th site that is initially prepared on the ground state $\left|\downarrow_{l}\right$>, it can be checked that after time $t$, the probability of transition to Rydberg state $\left|\uparrow_{l}\right$> is
@@
Figure 1: (a) Schematic illustration of the 1D Rydberg lattice under periodic driving field $\mathit{\Omega}\left(t\right)=\mathit{\Omega}_{0}\cos\left(\omega t\right)$. There are long-range interactions $V_{i,j}=\mathcal{V}_{0}/\left|i-j\right|^{6}$ between atoms in the Rydberg excited states $\left|\uparrow\right$>, and $\left|\downarrow\right$> represents the ground state of an atom. The bottom panels present the numerical results of time evolutions of the Rydberg density $\left<\hat{n}_{j}\right$> for different initial excitations and driving frequencies of driving fields. (a) Time evolutions of Rydberg density for simple spin cluster initial state under the resonant frequency $\omega=5.0867$, which is resonant with the Rydberg interaction and the system exhibits expansion dynamics of spin cluster. (b) Time evolutions of Rydberg density for a Gaussian distribution of spin cluster as initial state near the resonant frequency with $\omega=5.2867$. In contrast, the system exhibits oscillation dynamics in this case. The size of the system is set as $N=20$, and the strength of the field and interaction are taken as $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$, respectively.
Second, the interaction between the atoms in Rydberg states can be replaced by the effective detuning $\mathit{\Delta}_{\mathrm{eff}}\sim\mathcal{V}_{0}$, under the mean-field approximation. Then, inside the spin cluster, each atom can be described by the Hamiltonian $\hat{H}_{j}(t)=\mathit{\Omega}\left(t\right)\hat{\sigma}_{j}^{x}/2+\mathit{\Delta}_{\mathrm{eff}}\hat{n}_{j}$. According to the time-dependent perturbation theory, for an atom at site $j$ that is initially prepared in the Rydberg state $\left|\uparrow_{j}\right$>, the probability of transition to the ground state $\left|\downarrow_{j}\right$> at time $t$ is approximately
@@
which is suppressed when $\mathcal{V}_{0}\approx\omega\gg\mathit{\Omega}_{0}$ (see APPENDIX A for the derivation). This achieves a similar effect to that under the facilitation conditions [Marcuzzi2017, Ostmann2019, Magoni2021, Magoni2024, Brady2025], where the detuning is set to cancel out the interaction between two adjacent atoms. In this parameter regime, it is interesting to further reveal the dynamic behavior of the domain walls of the spin cluster.
In order to gain some intuitions and preliminary conclusions about the spin cluster dynamics, we perform numerical computations of time evolutions, observing the expansion dynamics of the spin cluster when the driving frequency is taken as $\omega=5.0867$, and a slight deviation of the driving frequency leads to oscillation dynamics of the spin cluster, in Figs. 1(b) and (c), respectively. In Fig. 1(b), we present numerical results of the Rydberg density $\left<\hat{n}_{j}\right>(t)=\left<\Psi\left(t\right)\right|\hat{n}_{j}\left|\Psi\left(t\right)\right$> as a function of time for the initial excitation in Eq. (2), that is $\left|\Psi\left(0\right)\right>=\left|\psi\left(j_{1,}j_{2}\right)\right$> with $j_{1}=9$ and $j_{2}=12$, for the first case, which is the simplest initial state to investigate the spin cluster dynamics and collision of domain walls of the driven Rydberg lattice. The evolved state
@@
is calculated under the time-dependent Hamiltonian in Eq. (1) using the fourth-order Runge-Kutta method with time steps of length $\Delta t=4\times 10^{-4}$ and the Hamiltonian is updated in each time step, where $\hat{\mathcal{T}}$ is the time-ordering operator and $\mathrm{i}=\sqrt{-1}$ (see APPENDIX B for more details). Note that the spin cluster state in Eq. (2) consists of all momentum components. Therefore, it is worth considering Gaussian distribution with central momentum $k_{0}$ to investigate the contribution of different $k$ components to spin cluster dynamics. In Fig. 1(c), we take a Gaussian distribution with width $d=2$ as initial state, which has the following form
$\left|\Psi_{\mathrm{Gau.}}\left(0\right)\right>=\Lambda^{-1}\sum_{j_{1}}\mathrm{e}^{-\frac{\left(c-c_{0}\right)^{2}}{4d^{2}}}\mathrm{e}^{-\mathrm{i}k_{0}c}\left|\psi\left(j_{1},j_{1}+r_{0}-1\right)\right$>, (6)
where $c=(j_{1}+j_{2})/2=j_{1}+(r_{0}-1)/2$ is the center of mass position of spin cluster (here $r_{0}=4$ is taken), $c_{0}=10.5$ and $k_{0}=\pi$ are, respectively, the average center of mass position and momentum, and $\Lambda$ is the normalization constant. We analyze the mechanism of these dynamics in the next section.
@@ III Effective Hamiltonian and resonant frequency
Figure 2: Numerical results of time average overlap $\overline{O(t)}$ between the evolved states computed from the Rydberg spin Hamiltonian and the effective Hamiltonian for different driven frequency $\omega$. The initial states are all taken as the spin cluster states in Eq. (2) with the positions of domain walls $j_{1}=8$ and $j_{2}=11$. The size of the system is set as $N=18$, and the strength of the field and interaction are taken as $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$, respectively. The results indicate that the effective Hamiltonian works well in the high frequency drive region.
Based on the above analyses, the effective Hamiltonian should works well in the high frequency drive region. To verify this point, we provide the comparison by the overlap between the evolved states computed from the Rydberg spin Hamiltonian and the effective Hamiltonian, which is defined as
$O\left(t\right)=\left|\langle\Psi\left(t\right)|\Psi_{\mathrm{eff}}\left(t\right)\rangle\right|^{2},$ (8)
where $\left|\Psi\left(t\right)\right>=\hat{\mathcal{T}}\exp\left[-\mathrm{i}\int_{0}^{t}\hat{\mathcal{H}}\left(t^{\prime}\right)\textrm{d}t^{\prime}\right]\left|\Psi\left(0\right)\right$> and $\left|\Psi_{\mathrm{eff}}\left(t\right)\right>=\hat{\mathcal{T}}\exp\left[-\mathrm{i}\int_{0}^{t}\hat{H}_{\mathrm{eff}}\left(t^{\prime}\right)\textrm{d}t^{\prime}\right]\left|\Psi\left(0\right)\right$>, and the initial states are taken as the spin cluster states in Eq. (2). In Fig. 2, we present the numerical results of time average overlap
$\overline{O(t)}=\frac{1}{t_{1}}\int_{0}^{t_{1}}O\left(t\right)\mathrm{d}t$ (9)
for different driven frequency $\omega$, and we take $t_{1}=10$. We can see that $\overline{O(t)}$ close to $1$ for high frequency drive $\omega\gg\mathit{\Omega}_{0}$, which verifies that the effective Hamiltonian works well in the high frequency drive region.
Figure 3: Time evolutions of Rydberg density $\left<\hat{n}_{j}\right$> and density variance $\delta\sigma(t)=\sigma(t)-\sigma(0)$ for the system under the resonant frequency of the driving field with different initial states. The frequency of the driving field is set as $\omega=5.0867$, which is resonant with the Rydberg interaction. The top panels present time evolution of Rydberg density $\left<\hat{n}_{j}\right$> for the initial excitations of (a1) simple spin cluster $\left|\psi\left(j_{1},j_{2}\right)\right$> in Eq. (2) with the positions of domain walls $j_{1}=46$ and $j_{2}=55$; the Gaussian distribution of the spin cluster in Eq. (6) with different central momenta (b1) $k_{0}=0$, (c1) $k_{0}=\pi/2$ and (d1) $k_{0}=\pi$. The width, central position of the Gaussian distribution and size of the spin cluster are taken as $d=4$, $c_{0}=50.5$ and $r_{0}=10$, respectively. The bottom panels (a2)-(d2) show the density variance $\delta\sigma(t)$ on double-logarithmic scales (black dots) corresponding to the Rydberg density in (a1)-(d1), and the solid lines represent the linear fittings for the data. The size of the system is set as $N=100$, and the strength of the field and interaction are taken as $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$, respectively. In the cases of (a)-(c), it can be observed that the exponent $\beta$ of density variances decrease due to the collision of the domain walls. While the case in (c) with central momentum $k_{0}=\pi$, the spin cluster exhibits ballistic expansion with exponent $\beta=2$.
@@
which is a constant for large $r$. This means that for the fixed $k$ and $t$, the Hamiltonian $\hat{H}_{k}(t)$ describes a hopping particle in a lattice with skew potential $\mathcal{F}r$, which is closely related to the Bloch oscillations [Bloch1929, Wannier1960].
In the large-$N$ limit and for a finite $\mathcal{F}$, the eigenstates of the Hamiltonian in Eq. (11) are all localized, forming the celebrated Wannier-Stark ladder states [Wannier1960]. We consider the eigenstates localized at $r\gg 1$, which have the form
@@
$\left|\Psi\left(0\right)\right>=\sum_{k}\mathcal{A}_{k}\left|\Psi_{k}\left(0\right)\right>\otimes\left|k\right$>, (14)
@@
and left-hand side multiplied by $\left<\phi_{m,k}\right|$, we obtain the equation of motion for amplitude $\mathcal{C}_{n,k}(t)$, that is
@@
Note, however, that $\left|\phi_{n,k}\right\rangle$ is time dependent. Submitting the expression of $\left|\phi_{n,k}\right\rangle$ [see Eq. (13)] into the above equation, and using the recurrence relation of Bessel function $2\partial_{x}J_{n}(x)=J_{n-1}(x)-J_{n+1}(x)$, we have
@@
Figure 4: Total Rydberg density $\mathcal{N}(t)$ as a function of time for the resonant cases in Fig. 3(a1)-(d1). The frequency of the driving field is set as the resonant value $\omega=5.0867$. The size of the system is set as $N=100$, and the strength of the field and interaction are taken as $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$, respectively. Combining with the results in Fig. 3, one can see that the total Rydberg density increase due to the collision of domain walls, and the total Rydberg density is conserved for the case with central momentum $k_{0}=\pi$.
Figure 5: Time evolution of Rydberg density $\left<\hat{n}_{j}\right$>, density variance $\delta\sigma(t)=\sigma(t)-\sigma(0)$ and autocorrelation function $\mathcal{A}(t)$ for the system near the resonant frequency of the driving field with different initial states. The frequency of the driving field is set as $\omega=5.1367$. The top panels (a1)-(d1) present time evolutions of Rydberg density $\left<\hat{n}_{j}\right$> for the initial excitations taken as the same as those in Fig. 3(a1)-(d1). The middle panels (a2)-(d2) show the density variance $\delta\sigma(t)$ corresponding to the Rydberg density in (a1)-(d1). The bottom panels show the autocorrelation function $\mathcal{A}(t)$ between the initial states and evolved states corresponding to the cases in (a1)-(d1). The size of the system is set as $N=100$, and the strength of the field and interaction are taken as $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$, respectively. The periodic boundary condition is taken for the system. For the cases in (a)-(c), the systems exhibit confinement dynamics of spin cluster without revival due to the near resonant condition and the collision of spin cluster domain walls, while for the Gaussian distribution with $k_{0}=\pi$ in (d), the spin cluster exhibits oscillation dynamics with high-amplitude revival due to the absent of the collision.